use laplace transfrom to identify F(x)

y''-y'-6y=0 given y(0)=1 y'(0)=2
Answer #1

Solution y-y-640 y 10) El y lol = 2 taking replace transform both-side of Lly-y-cy) = LIO 11y) - Lig) - 644) = 0 take Lly{ 6 -3) (842) 1-7 J() 어 6 19 (s)) (+)) 글 (+) (( T h ។ 세 오 t To 十 8-3 S ( 31 니e + c :) (re T (KOR (4) ) bts beldi sinec